# When do World Models Successfully Learn Dynamical Systems?

### v1

@@ A.6 Further Dataset Details
A.6.1 Initial Conditions
where $x,x^{\prime}\in\mathbb{T}^{2}$, $m>0$ is the inverse correlation length scale and $\nu>0$ is the smoothness parameter. Table 1 summarizes the values of $(\sigma,m,\nu)$ used in the datasets. The strength of the noise depends linearly on $\sigma$.
[…]
=1.2mm Parameter Value Noise type Normal Noise strength $\sigma$ $10$ Inverse correlation length scale $m$ in lattice step units $0.1$ Smoothness of Matern function $\nu$ $1$
A.6.2 Heat and Wave Equations
These datasets are generated on the flat torus $\mathbb{T}^{2}=[-1,1]^{2}$ by employing index wrapping i.e. circular boundary conditions. For all linear datasets, the data is linearly normalised to the range $[-1,1]$. The seeds used to generate $a$ for the operator $\nabla(a\nabla)$ can be found in the supplementary material, published with the paper. As noted in A.2, observability is a generic condition, and so we have observability for our modified operator almost surely.
@@ A.6.2 Heat and Wave Equations
A.6.3 KS Equation
[…]
=1.2mm Name KSE (low-res) KSE (super-res) Resolution $64\times 64$ $256\times 256$ Compression factor $16$ $16$ History size $16$ $7$ Frames (per init. cond.) $2000$ $2000$ No. of init. cond. (train) $28$ $28$ No. of init. cond. (test) $5$ $5$ Frame skip $10$ $10$ Burn-in frames $500$ $500$ $\Delta t$ $0.01$ $0.01$
@@ A.6.3 KS Equation
A.7 Setup and Configuration of Neural Network Training
A.7.1 Linear Equations
Here, a single autoregressive linear layer is sufficient, with input and output sizes matching the relevant dataset for each model. This has the added benefit of entirely preventing over-fitting, since the number of degrees of freedom in the model matches exactly that of the reconstruction map, when the history size is set to the compression factor, assuming there is enough data to properly constrain the model. We used the Adam optimiser [54], and an ablation study was conducted to verify the optimality of the hyperparameters, albeit with a coarse grid since the purpose of these examples is to empirically substantiate the learning of the time evolution operator and confirm the theory, and each run represents significant computational expense A.8. A learning rate of $1\times 10^{-5}$ was used, and the number of steps given in Section A.8. We employed Mean Squared Error (MSE) as the objective function during training. Such a low learning rate and high number of steps is required due to the ill-conditioned nature of these problems.
A.7.2 KS equation
[…]
=1.2mm Parameter Value Features per block (Low-res) $[64,128,256,512]$ Features per block (Super-res) $[64,128,256]$ Learning rate $0.0005$ Batch size $768$
@@ A.7.2 KS equation
[…]
A.7.3 Flow around a cylinder
NVIDIA’s LongVideoGAN [48] is designed to generate realistic video sequences with temporal consistency. Its architecture comprises two key components: a low-resolution GAN that captures time dynamics and a high-resolution GAN that enhances vortex detail by upscaling turbulent flows synthesised by the low-resolution GAN. Together, they ensure both the plausibility of individual frames and coherent motion throughout the video. The low-resolution GAN generates video sequences with resolution $64\times 36$ pixels. It starts by sampling an 8-dimensional latent vector for each frame from a standard Gaussian distribution. To enforce temporal consistency, these vectors are smoothed using Kaiser low-pass filters [69], which reduce high-frequency noise while preserving long-term patterns. Each value in the filtered sequence is influenced by the entire sequence, including future frames. After smoothing, the vectors are processed through fully connected and 3D convolutional layers to produce low-resolution frames that capture the essential dynamics and structure of the video. The high-resolution GAN upsamples the low-resolution video to resolutions such as $1024\times 576$ pixels, using the latter as conditional input. To generate the $n$-th high-resolution frame, it conditions on frames $n-4$ to $n+4$ from the low-resolution sequence, preserving temporal consistency. Aside from this conditioning, its architecture closely follows that of StyleGAN3 [70], incorporating convolutional and upsampling layers. During training, it is conditioned on real low-resolution videos aligned with the target high-resolution outputs. DiffAug [71] is applied to improve robustness through augmentations such as translation and cutout, while R1 regularisation [72] stabilises training by penalising the discriminator only on real samples. Once trained, the low-resolution GAN generates a video, which is then refined and upscaled by the high-resolution GAN to produce the final output.
@@ A.7.3 Flow around a cylinder
To assess how well VideoGAN and the KSE model capture turbulent flow dynamics, we use a temporal correlation metric. For a video of length $T$, we compute the Pearson correlation $\rho(\Delta t)$ between the values of a specific pixel at time $t$ and $t+\Delta t$ for all $t\in 1,\dots,T-\Delta t$. This measures how predictive a pixel’s value is over time. High $\rho(\Delta t)$ indicates strong temporal consistency. Additional evaluation metrics specific to VideoGAN are provided in Appendix A.10.
@@
NVIDIA’s LongVideoGAN requires a 16!:!916!:!916 ! : ! 9 image format. To reshape the data to $1024\times 576$ resolution, we added $24$ white columns to the left side of the image and removed $12$ rows from both the top and bottom. These images were then downscaled to $64\times 36$ using Python’s Pillow library for low-resolution input. Grayscale images were converted to RGB by replicating the single channel three times. The resulting $19,990$ images were split into $11$ directories, each forming a non-overlapping video of approximately $1,817$ frames for training. Both GANs were trained on 2 NVIDIA A100 GPUs with a batch size of 8 and gradient accumulation of 1. All other hyperparameters followed those from the original LongVideoGAN paper and code.
A.8 Computational Cost
A.8.1 Linear Equations
@@ A.8.2 Non-linear Equation
A.8.3 Large Eddy Simulation
The LES was run on Intel Xeon "Skylake" Gold 6132 CPUs running at 2.6 GHz and equipped with 96 GB of RAM. The simulation used 20 nodes, each with 28 cores, and ran for approximately 20 days, for a total of 1,440 core weeks [10].
A.8.4 LongVideoGAN
Our training of the Video GAN using $2$ A100 GPUs took about $7$ days, and the generation of a $1000$ frame video takes about $32$ seconds using one NVIDIA R6000. (On the Anonymous University HPC cluster it then takes another $60$ seconds to save a video of length $1000$.) The generation process is GPU RAM intensive and will require more than one GPU to generate videos with more than $10,000$ frames and/or higher resolution. We note that GANs are particularly fast for this kind of dataset, see [73] for more details. At 20 days for 100,000 frames, the CFD simulation produces 1 frame approximately every $17s$. Our model produces 1 frame every $0.032s$, which is a speed up factor of over 500x. We emphasise that this speed up is precisely the point of these experiments, and CFD simulations remain prohibitively expensive.
A.9 Dataset Snapshots
@@
A.10 Further Evaluations
A.10.1 Least-Squares Regression
@@
Figure 14: Average $L_{1}$ error as a function of the model history size for the least-squares regression case
Figure 15: $L_{\infty}$ error as a function of the model history size for the least-squares regression case
A.10.2 VideoGAN Generalization
@@
$V$ |$|\{g\in G:V=\text{argmin}_{v\in\mathcal{V}}d(g,v)\}|$
@@
Table 5: (a) The minimum Euclidean distance measured from each of the $4$ video sets over all $g\in G$ where $G$ is the set of the $50$ clips used for this evaluation. (b) Counts of how often each of the four sets is the closest to a generated clip, where $\mathcal{V}=\{V^{\text{Train}},V^{\text{Test1}},V^{\text{Test2}},V^{\text{Test3}}\}$.
@@
A.10.3 Further Temporal Correlations
@@
Figure 17: Temporal correlations of generated videos of length $301$ (left) and $1000$ (right) for pixels $1$, $2$, $3$ and $4$ (top to bottom).
@@
A.11 Navier Stokes Equations for Large Eddy Simulation (LES)
In section 3.3, we introduced the Navier–Stokes equations, on which turbulent flows rely. More specific we obtain our data by Large Eddy Simulation (LES) which is based on the concept of scale separation in turbulent flows, allowing for the large, energy-containing eddies to be resolved explicitly while modelling the effects of the smaller, subgrid-scale motions [74]. This is achieved through a spatial filtering operation, which separates the flow $\varphi(x,t)$ into a filtered (resolved) component $\overline{\varphi}(x,t)$ and a subgrid-scale component $\dot{\varphi}(x,t)$, such that
@@ A.11 Navier Stokes Equations for Large Eddy Simulation (LES)
[…]
@@
Applying this filtering operation to the NSE, where it is assumed [74] that the filtering and differential operators commute, yields the filtered LES form:

### v2

@@ Appendix G Further Dataset Details
G.1 Linear Equations
The heat and wave equation solutions are constructed by forward Euler, an explicit method. This scheme, while simple, is known to converge for appropriately small time steps for these equations (heat_wave_convergence). Since the Laplacian $\nabla^{2}$ is not observable under the tokenization operator (Theorem 6) we generate a modified Laplacian $\nabla(a\nabla)$ where $a$ is a stochastically generated positive definite diagonal matrix. Examples of the low-res (super-res) heat (wave) dataset can be found in the left hand column of Fig. 10 (Fig. 11), in the Appendix. The compression factor was selected to roughly match that of the LongVideoGAN model. Table 4 gives further details on the production of these datasets.
These datasets are generated on the flat torus $\mathbb{T}^{2}=[-1,1]^{2}$ by employing index wrapping i.e. circular boundary conditions. For all linear datasets, the data is linearly normalised to the range [1,1][-1,1]. The seeds used to generate $a$ for the operator $\nabla(a\nabla)$ can be found in the supplementary material, published with the paper. As noted in B, observability is a generic condition, and so we have observability for our modified operator almost surely.
@@ G.1 Linear Equations
G.2 KS Equation
A large amount of data was created by adopting the ETDRK4 (an exponential 4th order Runge-Kutta method) (etdrk4), which we modified to produce solutions to the 2D equation. Notably, the 0th mode in Fourier space was zeroed out to prevent the growth of the mean $\bar{u}$ (see Eq. 2.2 in 2dkse_thesis). The precise numerical parameters used can be found in Table 5, and still frames from the super-res dataset can be found in the left column in Fig. 9.
Name KSE (low-res) KSE (super-res)
Resolution $64\times 64$ $256\times 256$
Compression factor $16$ $16$
History size $16$ 77
Frames (per init. cond.) $2000$ $2000$
No. of init. cond. (train) $28$ $28$
No. of init. cond. (test) 55 55
Frame skip $10$ $10$
Burn-in frames $500$ $500$
$\Delta t$ $0.01$ $0.01$
@@ G.2 KS Equation
G.3 Initial Conditions
where $x,x^{\prime}\in\mathbb{T}^{2}$, $m>0$ is the inverse correlation length scale and $\nu>0$ is the smoothness parameter. Table 6 summarizes the values of $(\sigma,m,\nu)$ used in the datasets. The strength of the noise depends linearly on $\sigma$. This is a standard model of non trivial spatial random structures that is widely used in e.g. geo-statistics or the theory of PDEs with random coefficients. The resulting Gaussian random field is then transformed with the exponential function to give positive values for the conductivity. This is implemented using the fast Fourier transform on the discrete torus: First FTT the noise, then multiply with the FFT of the square root of the Matern covariance operator, then apply the inverse FFT, and then exponentiate.
Parameter Value
Noise type Normal
Noise strength $\sigma$ $10$
Inverse correlation length scale $m$ in lattice step units $0.1$
Smoothness of Matern function $\nu$ 11
G.4 Flow Around a Cylinder
For the experiments involving flow around a cylinder at Reynolds number 3900, we used a publicly available dataset winhart2024data provided by the authors of drygala2022generative. The dataset consists of 100,000 greyscale images with a resolution of $1000\times 600$ pixels, capturing the characteristic Kármán vortex street - a coherent structure of alternating vortices aligned with the cylinder’s axis. The data was generated using Large Eddy Simulation (LES), and the resulting unsteady velocity fields were processed via a projection mapping technique that preserves ergodicity (ergodicity_basics) in a reduced state space. For details on the LES numerical setup, we refer to drygala2022generative. The details on the setup and configuration of neural network training can be found in Section H.
Appendix H Setup and Configuration of Neural Network Training
H.1 Linear Equations
Here, a single autoregressive linear layer is sufficient, with input and output sizes matching the relevant dataset for each model. This has the added benefit of entirely preventing over-fitting, since the number of degrees of freedom in the model matches exactly that of the reconstruction map, when the history size is set to the compression factor, assuming there is enough data to properly constrain the model. We used the Adam optimiser (adam), and an ablation study was conducted to verify the optimality of the hyperparameters, albeit with a coarse grid since the purpose of these examples is to empirically substantiate the learning of the time evolution operator and confirm the theory, and each run represents significant computational expense (I). A learning rate of $1\times 10^{-5}$ was used, and the number of steps given in Section I. We employed Mean Squared Error (MSE) as the objective function during training. Such a low learning rate and high number of steps is required due to the ill-conditioned nature of these problems.
H.2 KS equation
Parameter Value
Features per block (Low-res) $[64,128,256,512]$
Features per block (Super-res) $[64,128,256]$
Learning rate $0.0005$
Batch size $768$
@@ H.2 KS equation
Figure 8: Architecture of the UNet in the low-res case, for the KS model. The very last layer is a final convolution to send the output from $\text{history size}\times\text{resolution}\rightarrow\text{resolution}$. All layers use a ReLu activation, and batch normalisation.
H.3 Flow around a cylinder
NVIDIA’s LongVideoGAN (Long_video_gan) is designed to generate realistic video sequences with temporal consistency. Its architecture comprises two key components: a low-resolution GAN that captures time dynamics and a high-resolution GAN that enhances vortex detail by upscaling turbulent flows synthesised by the low-resolution GAN. Together, they ensure both the plausibility of individual frames and coherent motion throughout the video. The low-resolution GAN generates video sequences with resolution $64\times 36$ pixels. It starts by sampling an 8-dimensional latent vector for each frame from a standard Gaussian distribution. To enforce temporal consistency, these vectors are smoothed using Kaiser low-pass filters (1974Nonrecursive), which reduce high-frequency noise while preserving long-term patterns. Each value in the filtered sequence is influenced by the entire sequence, including future frames. After smoothing, the vectors are processed through fully connected and 3D convolutional layers to produce low-resolution frames that capture the essential dynamics and structure of the video. The high-resolution GAN upsamples the low-resolution video to resolutions such as $1024\times 576$ pixels, using the latter as conditional input. To generate the $n$-th high-resolution frame, it conditions on frames n4n-4 to $n+4$ from the low-resolution sequence, preserving temporal consistency. Aside from this conditioning, its architecture closely follows that of StyleGAN3 (Karras2021), incorporating convolutional and upsampling layers. During training, it is conditioned on real low-resolution videos aligned with the target high-resolution outputs. DiffAug (zhao2020differentiable) is applied to improve robustness through augmentations such as translation and cutout, while R1 regularisation (r1_reg) stabilises training by penalising the discriminator only on real samples. Once trained, the low-resolution GAN generates a video, which is then refined and upscaled by the high-resolution GAN to produce the final output.
@@ H.3 Flow around a cylinder
To assess how well VideoGAN and the KSE model capture turbulent flow dynamics, we use a temporal correlation metric. For a video of length $T$, we compute the Pearson correlation $\rho(\Delta t)$ between the values of a specific pixel at time $t$ and $t+\Delta t$ for all $t\in 1,\dots,T-\Delta t$. This measures how predictive a pixel’s value is over time. High $\rho(\Delta t)$ indicates strong temporal consistency. Additional evaluation metrics specific to VideoGAN are provided in Appendix K.
@@
NVIDIA’s LongVideoGAN requires a $16:9$ image format. To reshape the data to $1024\times 576$ resolution, we added $24$ white columns to the left side of the image and removed $12$ rows from both the top and bottom. These images were then downscaled to $64\times 36$ using Python’s Pillow library for low-resolution input. Grayscale images were converted to RGB by replicating the single channel three times. The resulting $19,990$ images were split into $11$ directories, each forming a non-overlapping video of approximately $1,817$ frames for training. Both GANs were trained on 2 NVIDIA A100 GPUs with a batch size of 8 and gradient accumulation of 1. All other hyperparameters followed those from the original LongVideoGAN paper and code.
Appendix I Computational Cost
I.1 Linear Equations
@@ I.2 Non-linear Equation
I.3 Large Eddy Simulation
The LES was run on Intel Xeon "Skylake" Gold 6132 CPUs running at 2.6 GHz and equipped with 96 GB of RAM. The simulation used 20 nodes, each with 28 cores, and ran for approximately 20 days, for a total of 1,440 core weeks (drygala2022generative).
I.4 LongVideoGAN
Our training of the Video GAN using $2$ A100 GPUs took about $7$ days, and the generation of a $1000$ frame video takes about $32$ seconds using one NVIDIA R6000. (On the Anonymous University HPC cluster it then takes another $60$ seconds to save a video of length $1000$.) The generation process is GPU RAM intensive and will require more than one GPU to generate videos with more than $10,000$ frames and/or higher resolution. We note that GANs are particularly fast for this kind of dataset, see drygalaross2024 for more details. At 20 days for 100,000 frames, the CFD simulation produces 1 frame approximately every $17s$. Our model produces 1 frame every $0.032s$, which is a speed up factor of over 500x. We emphasise that this speed up is precisely the point of these experiments, and CFD simulations remain prohibitively expensive.
Appendix J Dataset Snapshots
@@ I.4 LongVideoGAN
Appendix K Further Evaluations
K.1 Least-Squares Regression
@@
Figure 13: Average $L_{1}$ error as a function of the model history size, where the model is ‘trained’ by least-squares regression
Figure 14: $L_{\infty}$ error as a function of the model history size, where the model is ‘trained’ by least-squares regression
K.2 VideoGAN Generalization
@@
$V$ $|\{g\in G:V=\text{argmin}_{v\in\mathcal{V}}d(g,v)\}|$
@@
Table 8: (a) The minimum Euclidean distance measured from each of the 44 video sets over all $g\in G$ where $G$ is the set of the $50$ clips used for this evaluation. (b) Counts of how often each of the four sets is the closest to a generated clip, where $\mathcal{V}=\{V^{\text{Train}},V^{\text{Test1}},V^{\text{Test2}},V^{\text{Test3}}\}$.
@@
K.3 Further Temporal Correlations
@@
Figure 16: Temporal correlations of generated videos of length $301$ (left) and $1000$ (right) for pixels 11, 22, 33 and 44 (top to bottom).
@@
Appendix L Navier Stokes Equations for Large Eddy Simulation (LES)
In section˜3.3, we introduced the Navier–Stokes equations, on which turbulent flows rely. More specific we obtain our data by Large Eddy Simulation (LES) which is based on the concept of scale separation in turbulent flows, allowing for the large, energy-containing eddies to be resolved explicitly while modelling the effects of the smaller, subgrid-scale motions (sagaut2005large). This is achieved through a spatial filtering operation, which separates the flow $\varphi(x,t)$ into a filtered (resolved) component $\overline{\varphi}(x,t)$ and a subgrid-scale component $\dot{\varphi}(x,t)$, such that
@@
$\overline{\varphi}(x,t)=\int\limits_{-\infty}^{\infty}\int\limits_{-\infty}^{\infty}G(x-x^{\prime},t-t^{\prime})\varphi(x^{\prime},t^{\prime})\,dt^{\prime}\,dx^{\prime},$ (24)
@@
Applying this filtering operation to the NSE, where it is assumed (sagaut2005large) that the filtering and differential operators commute, yields the filtered LES form: